If x is 125% of y. y is what % of X.
x=125%y
y=X/125%

Answer by ptaylor(2198)

If x is 125% of y. y is what % of X.
x=125%y
y=X/125%
YOU ARE RIGHT ON!!! YOU JUST HAVE TO FINISH IT.
Change 125% to a fraction or decimal
125%=`125/100=1.25
y=x/1.25=0.8x (1 divided by 1.25) now change back to a percent(0.8=80/100=80%)
y=80% of x
Hope this helps---ptaylor